# Order Parameter Discovery for Quantum Many-Body Systems

### v1

@@ V.2 Order parameters discovery
Next, we experimentally demonstrate the validity of the method used for order parameter discovery (Section IV), and in doing so, propose a method for understanding the structure of the optimal observable. We proceed with the ANNNI model (16) by considering the phase diagram in the region $(\kappa\quad h)\in\mathcal{R}=[0.5,2.1]\times[0,1.6]$ and begin with the use of the RDMs of single spin sites. Let $M$ denote the optimal observable for the order parameter discovery problem in (IV), where the details concerning its solution are expanded in Appendix B. Optimizing for a single site observable, we find that $M\approx\mathds{I}-\sigma_{x}$, which can be interpreted as magnetization.
Although this observable could serve as an order parameter for the Ising-like transition, marked in green in Figure 4, it fails to detect the other transitions in the ANNNI model, namely the KT and PT transitions, which are highlighted in orange and blue, respectively, in Figure 4. Consequently, to identify all the phase transitions, one must refer to Section IV and instead develop a multi-site observable that can capture all the transitions.
@@
In Figure 5 (a) we present the expectations of the observable $M$ applied to the RDMs $\rho_{0}(\kappa_{i},h_{i})$, for a finite lattice of parameters $\{(\kappa_{i},h_{i})\}$, following the optimization process of obtaining a relevant order parameter $M$ for this phase transition. We note that in the optimization process, the entire region, $(\kappa\quad h)\in\mathcal{R}=[0.0,2.1]\times[0,1.6]$, was used, which included all phase transitions present. Various values of $\gamma$ in (IV) ranging from $[1,100]$ were also used, and it was found to have little effect in detecting each phase transition.
@@
with |$\ket{\Psi^{+}}=(\ket{\uparrow\downarrow}+\ket{\downarrow\uparrow})/\sqrt{2}$ (Bell’s state), so the ordered phase for $M^{(4)}$ is the antiphase. This can be justified by the modulated structure of the antiphase $|\ket{\cdots\leftarrow\leftarrow\rightarrow\rightarrow\cdots}$ (see Appendix F for a comprehensive introduction to ANNNI). Indeed we see that $\operatorname{Tr}\left(M^{(4)}\ket{\varphi\varphi}\bra{\varphi\varphi}\right)\neq 0$ for |$\ket{\varphi}=\ket{\leftarrow}$ and |$\ket{\varphi}=\ket{\rightarrow}$, whereas $\operatorname{Tr}\left(M^{(4)}\ket{\uparrow\uparrow}\bra{\uparrow\uparrow}\right)=0$ (paramagnetic).
@@ V.4 The cluster Hamiltonian
After constructing the phase diagram for the ANNNI model, we will now showcase the versatility of the fidelity vector field method by applying it to other models. Specifically, we apply the method described in Section III to the cluster Hamiltonian, defined by the Hamiltonian:
@@
When the two parameters of the Hamiltonian are equal we expect a phase transition between a trivial phase ($h>K$) and the Symmetry Protected Topological (SPT) phase ($h<K$). To obtain the ground states of the Hamiltonian, we use DMRG, as described in Appendix D and we repeat the calculations as outlined in subsection V.1 to construct the phase diagram for the cluster Hamiltonian, of which is shown in Figure 9.
@@
To conclude this section, we have demonstrated that the RFS vector field accurately reproduces the phase diagrams of both the ANNNI and cluster Hamiltonians, showcasing its robustness and reliability in capturing the phase transitions of diverse systems.
@@ Appendix A Construction of the phase diagram
for a fixed step $h$ (which defines the lattice of parameters) and $d_{i},d_{j}\in\{-1,0,1\}$. We can immediately verify the computational advantage of the approach since adjacent lattice points share a fidelity perturbation. For example $f(\boldsymbol{\lambda}_{i,j},h\mathbf{e}_{1})=f(\boldsymbol{\lambda}_{i,j+1},-h\mathbf{e}_{1})$. In Figure 10, the fidelity perturbations are represented with the symbol and placed between adjacent lattice points to emphasize the concept of sharing. By considering the finite differences approximation of the second derivative777 For a function $f:\mathbb{R}\to\mathbb{R}$, under sufficient smoothing conditions, we make use of the following approximation $\frac{\mathrm{d}^{2}f}{\mathrm{d}x^{2}}\approx$ $\frac{f(x+h)+f(x-h)-2f(x)}{h^{2}}$ (24) for some $h>0$. , and by noting that $f(\boldsymbol{\lambda}_{i,j},\mathbf{0})=1$ (for all valid $i,j$), we obtain
$\tilde{g}(\boldsymbol{\lambda}_{i,j})=$ $\frac{4-\sum_{\boldsymbol{\delta}\in\Omega}f(\boldsymbol{\lambda}_{i,j},\boldsymbol{\delta})}{h^{2}}$ (25)
@@ Appendix B Solution of the order parameter discovery problem
In this section, we expand on the solution of the order parameter discovery problem defined in (IV). We restate the latter for clarity and convenience, so
$\underset{M\in\mathcal{S}_{m}}{\min}$ $-\frac{1}{|I^{+}|}\sum_{i\in I^{+}}\langle M\rangle_{i}^{2}+\frac{\gamma}{|I^{-}|}\sum_{j\in I^{-}}\langle M\rangle_{j}^{2},$
@@
with $\langle M\rangle_{i}\coloneqq\operatorname{Tr}(\rho_{0}(\boldsymbol{\lambda}_{i})M)$.
@@
Assuming $M\in\mathcal{S}_{m}$, for some finite index set $I$, we note that
$\sum_{i\in I}\langle M\rangle_{i}^{2}=$ $\sum_{i\in I}\left(\operatorname{Tr}(\rho_{0}(\boldsymbol{\lambda}_{i})M)\right)^{2}$ (34a)
${=}$ $\sum_{i\in I}\left(\operatorname{\mathrm{vec}_{r}}(M)^{\dagger}\operatorname{\mathrm{vec}_{r}}(\rho_{i})\right)^{2}$ (34b)
$=$ $\mathbf{x}^{\dagger}\left(\sum_{i\in I}\mathbf{r}_{i}\mathbf{r}_{i}^{\dagger}\right)\mathbf{x},$ (34c)
with $\mathbf{x}\coloneqq\operatorname{\mathrm{vec}_{r}}(M)$ and $\mathbf{r}_{i}\coloneqq\operatorname{\mathrm{vec}_{r}}(\rho_{0}(\boldsymbol{\lambda}_{i}))$. We note that the vectors $\mathbf{x}$ and $\mathbf{r}_{i}$ belong to $\widehat{\mathcal{S}_{m}}$.
We use the result in (34a) to rewrite the optimization problem in (IV) as
@@
$A\coloneqq$ $-\frac{1}{|I^{+}|}\sum_{i\in I^{+}}\mathbf{r}_{i}\mathbf{r}_{i}^{\dagger}+\frac{\gamma}{|I^{-}|}\sum_{j\in I^{-}}\mathbf{r}_{j}\mathbf{r}_{j}^{\dagger}.$ (36)
The matrix $A$ (determined by data) is Hermitian, so the objective in (35a) is real and well-defined (even in the absence of constraint (35c)). We expand on the condition for the non-degeneracy of the problem introduced in Section IV. We impose that the matrix $A$ is indefinite [42], where the Hermitian structure is fulfilled by the construction in (36). In other words, we require the matrix $A$ to have both (strictly) positive and negative eigenvalues. The construction of $A$ shows the difference between two PSDs, so the tradeoff $\gamma>0$ is a sensible parameter since it can render the problem (numerically) degenerate.
@@
The optimization problem is non-convex since we do not assume $A\succcurlyeq 0$, indeed $A$ is required indefinite. However, as anticipated in Section IV, this optimization problem can be solved efficiently even in the case of the non-convexity of the objective (i.e. matrix $A$ is not PSD). Moreover, this is an exceptional case where strong duality888 Strong duality is equivalent to the duality gap is zero, that is the difference between primal and dual solutions.  [25] holds, provided that Slater’s constraint qualification is fulfilled. That is, there exists an $\mathbf{x}\in\widehat{\mathcal{S}_{m}}$ such that the inequality constraint (35b) holds strictly (i.e. not tight). In our case, an example is $\mathbf{x}=\operatorname{\mathrm{vec}_{r}}(\mathbb{I}_{m})/\sqrt{m+\epsilon}$ for any $\epsilon>0$, so $\|x\|^{2}_{2}<1$.
@@
then the constraint is fulfilled when $\alpha\geq-\lambda_{\mathrm{min}}(A)$, so the dual optimal is $\alpha^{\star}=-\lambda_{\mathrm{min}}(A)>0$. The space of the primal solutions is given by the eigenspace corresponding to the minimum eigenvalue of $A-\lambda_{\mathrm{min}}(A)\mathbb{I}_{m^{2}}$, that is the null space of the latter. Given any matrix $A$, we denote the null space of $A$ (i.e. the set of solutions of the homogeneous equation $Ax=0$) by $\mathrm{Null}(A)$. Let
@@
with $\|x^{\star}\|^{2}_{2}=1$, be an optimal solution999 We note that the vectorization operator $\operatorname{\mathrm{vec}_{r}}(\cdot):\mathbb{C}^{m\times m}\to\mathbb{C}^{m^{2}}$ is an isomorphism, so the expression $\mathbf{x}^{\star}=\operatorname{\mathrm{vec}_{r}}(M)$ implicitly means that the matrix $M$ is obtained uniquely from $\mathbf{x}^{\star}$ using the inverse of $\operatorname{\mathrm{vec}_{r}}$. , we show that if $\mathbf{x}^{\star}=\operatorname{\mathrm{vec}_{r}}(M)$, then $M=M^{\dagger}$. That is constraint (35c) is implied by the structure of matrix $A$.

### v3

@@ V.2 Order parameters discovery
Next, we experimentally demonstrate the validity of the method used for order parameter discovery (Section IV), and in doing so, propose a method for understanding the structure of the optimal observable. We proceed with the ANNNI model (19) by considering the phase diagram in the region $(\kappa\quad h)\in\mathcal{R}=[0.5,2.1]\times[0,1.6]$ and begin with the use of the RDMs of single spin sites. Let $M$ denote the optimal observable for the order parameter discovery problem in (IV), where the details concerning its solution are expanded in Appendix B. Optimizing for a single site observable, we find that $M\approx\mathds{I}-\sigma_{x}$, which can be interpreted as magnetization. While this observable could be used as an order parameter for the Ising-like transition, highlighted in green in Figure 4, it failed to detect other transitions present in the ANNNI model, including the KT and PT transitions, highlighted orange and blue in Figure 4 respectively. As a result, to capture more transitions one must make use of Section IV and instead uncover a multi-site observable to detect all phase transitions present.
@@
In Figure 5 (a) we present the expectations of the observable $M$ applied to the RDMs $\rho_{0}(\kappa_{i},h_{i})$, for a finite lattice of parameters $\{(\kappa_{i},h_{i})\}$, following the optimization process of obtaining a relevant order parameter $M$ for this phase transition. We note that in the optimization process, the entire region, $(\kappa\quad h)\in\mathcal{R}=[0.0,2.1]\times[0,1.6]$, was used, which included all phase transitions present.
@@
with |$\ket{\Psi^{+}}=(\ket{\uparrow\downarrow}+\ket{\downarrow\uparrow})/\sqrt{2}$ (Bell’s state), so the ordered phase for $M^{(4)}$ is the antiphase. This can be justified by the modulated structure of the antiphase $|\ket{\cdots\leftarrow\leftarrow\rightarrow\rightarrow\cdots}$ (see Appendix G for a comprehensive introduction to ANNNI). Indeed we see that $\operatorname{Tr}\left(M^{(4)}\ket{\varphi\varphi}\bra{\varphi\varphi}\right)\neq 0$ for |$\ket{\varphi}=\ket{\leftarrow}$ and |$\ket{\varphi}=\ket{\rightarrow}$, whereas $\operatorname{Tr}\left(M^{(4)}\ket{\uparrow\uparrow}\bra{\uparrow\uparrow}\right)=0$ (paramagnetic).
@@ V.3 Finite Size Scaling
To confirm that we belong to the Ising universality class, we fit the magnitude of the average observable near the critical point $h_{c}=1.0$. In this regime, the critical magnetization should follow the relationship $M\propto|L|^{\beta/\nu}$. Assuming $\nu=1.0$, we fit the values of the observed data and find $\beta=0.123$, which is close to the expected value of $\beta=\frac{1}{8}$ for the Ising universality class.
@@ V.4 The cluster Hamiltonian
Following the construction of the phase diagram for the ANNNI model, we next demonstrate the versatility of the fidelity vector field method by applying it to other models. Specifically, we apply the method described in Section III to the cluster Hamiltonian, defined by the Hamiltonian:
@@
When the two parameters of the Hamiltonian are equal, a phase transition is expected between a trivial phase (for $h>K$) and a Symmetry Protected Topological (SPT) phase (for $h<K$). To distinguish between these two phases, string order parameters (SOPs) are commonly used, as they can reveal hidden orders or symmetries within the system. In the case of the cluster Hamiltonian, the system exhibits certain symmetries, denoted by $G_{odd}=\prod_{k=0}Z_{2k+1}$ and $G_{even}=\prod_{k=0}Z_{2k}$.
From [38, 39], a general SOP can be expressed as
$S_{\Sigma}^{O^{L},O^{R}}=\lim_{|j-k|\to\infty}\left\langle\psi_{0}\left|O^{L}(j)\left(\prod_{i=j+1}^{k-1}\Sigma_{i}\right)O^{R}(k)\right|\psi_{0}\right\rangle$ (25)
where $\psi_{0}$ is the state of the system, $\prod_{i=j+1}^{k-1}\Sigma_{i}$ (in our case $G_{odd}$ or $G_{even}$) preserves the local symmetries of the system, and $O^{L/R}$ are appropriately chosen to detect the phase transition. In this particular system, a typical order parameter used for the phase transition is $\sigma^{x}\otimes\sigma^{z}\otimes I\otimes\sigma^{z}\otimes\sigma^{x}$, or another commonly used one is $\sigma^{x}\otimes\sigma^{y}\otimes\sigma^{z}\otimes\sigma^{y}\otimes\sigma^{z}$ as discussed in [40].
We begin by focusing on reconstructing the phase diagram as discussed in section Section III. To determine the ground states of the Hamiltonian, we employ DMRG, following the procedure described in Appendix E. We replicate the calculations outlined in subsection V.1 to construct the phase diagram for the cluster Hamiltonian, which is presented in Figure 9. The phase transition is visible, with a distinct boundary separating the two phases.
@@
Next, we turn our attention to identifying an order parameter for this phase transition by following the method described in Section IV. Using a five-site RDM, we derive an order parameter that effectively distinguishes between the different phases, as shown in 10.
Figure 10: Expectation of the optimal observable $M$ for cluster Hamiltonian. Here a chain length of $L=50$ spin sites was used and a five-site RDM was used when calculating the gradient of the reduced fidelity susceptibility.
Although this order parameter is a linear combination of various tensor products of Pauli operators, the two terms with the largest coefficients are $\sigma^{x}\otimes\sigma^{z}\otimes I\otimes\sigma^{z}\otimes\sigma^{x}$ and $\sigma^{x}\otimes\sigma^{y}\otimes\sigma^{z}\otimes\sigma^{y}\otimes\sigma^{z}$, both observables that reflect the symmetry of the SPT phase. This suggests that our observable has correctly identified the appropriate SOP and effectively exploited the system’s symmetries to distinguish between the two phases.
It is worth noting that similar calculations were performed using different sizes of RDMs, specifically one and two sites. These attempts, however, failed to detect the phase transition, indicating that these smaller RDMs do not capture enough information about the phases on their own.
@@
To conclude, we have demonstrated that the RFS vector field accurately reproduces the phase diagrams of both the ANNNI and cluster Hamiltonians, showcasing its robustness and reliability in capturing the phase transitions of diverse systems. Furthermore, we successfully identified an order parameter that leverages each system’s symmetries to distinguish between different phases, further validating our approach.
V.5 The Rydberg model spin chain
Figure 11: Schematic of the phase diagram for the Rydberg Model as a function of the parameters $\frac{\Delta}{\Omega}$ and $\left(\frac{R_{b}}{a|i-j|}\right)^{6}$ (see the main text for more details). The orange regions represent the emergence of various crystalline quantum phases with $\mathbb{Z}_{N}$ order.
V.5.1 Model description
The Rydberg model is a quantum many-body system that has attracted considerable interest, particularly its role in experimental quantum simulation platforms [41, 42, 43, 44]. The system is composed of a chain of atoms such that each atom can either remain in its ground state or be excited to a highly energetic Rydberg state, characterized by an electron occupying a very high orbital [45]. The model captures the dynamics of strongly interacting systems, as atoms in Rydberg states exhibit long-range dipole-dipole interactions, significantly affecting the system’s collective behavior. Using these interactions, a variety of exotic quantum phases can be simulated, including quantum phase transitions and spin-liquid phases, making the Rydberg model an essential tool for probing non-equilibrium phenomena in quantum simulators [46].
Figure 12: Phase diagram for the Rydberg Model obtained with a one-site RDM from a chain of length $L=50$, with long-range interactions truncated at the fourth nearest neighbor. The lower crystalline phase corresponds to the $\mathbb{Z}_{2}$ crystalline phase, while the upper corresponds to the $\mathbb{Z}_{3}$ crystalline phase.
Mathematically, the Rydberg model can be formulated as a spin-$1/2$ system, where each site of the chain represents an atom that can either be in the ground state (represented by the spin state $\left|0\right\rangle$) or in the Rydberg state (represented by the spin state $\left|1\right\rangle$). The Hamiltonian of the Rydberg model is typically expressed as
$H=\Omega\sum_{i}\sigma_{i}^{x}-\Delta\sum_{i}n_{i}+\sum_{i<j}V_{ij}n_{i}n_{j},$ (26)
where $\sigma_{i}^{x}$ is the Pauli-x matrix acting on-site $i$, and $n_{i}=\left|1\right\rangle\left\langle 1\right|_{i}$ is the number operator that projects onto the Rydberg state at site $i$. The parameters $\Omega$ and $\Delta$ represent the Rabi frequency and detuning, respectively, which control the driving field and energy offset of the system. The second term in the Hamiltonian describes the interaction between atoms in their Rydberg states, with $V_{ij}(r)\propto 1/r^{6}$ representing the strength of the interaction between atoms at sites $i$ and $j$. For $\Delta>0$, various spatially ordered phases emerge due to the competition between two effects: the detuning term, which favors a high population of atoms in the Rydberg state, and the Rydberg blockade, which prevents the simultaneous excitation of atoms that are closer together than the blockade radius $R_{b}$, where the interaction strength $V(R_{b})$ is equal to the Rabi frequency $\Omega$.
We define the distance between adjacent spins as $a$, thus $r=|i-j|a$, and we also define the Rydberg Radius as $V(R_{b})=\Omega$. Thus the Hamiltonian can be rewritten accordingly;
$H=\sum_{i}\sigma_{i}^{x}-\frac{\Delta}{\Omega}\sum_{i}n_{i}+\left(\frac{R_{b}}{a}\right)^{6}\sum_{i<j}\frac{1}{|i-j|^{6}}n_{i}n_{j},$ (27)
It is the parameters $\frac{\Delta}{\Omega}$ and $\left(\frac{R_{b}}{a|i-j|}\right)^{6}$ that we will vary in the following section.
In the limit of strong interactions, the Rydberg blockade phenomenon becomes significant, where two neighboring atoms cannot both be excited to the Rydberg state due to the large energy cost. This constraint leads to the emergence of various quantum phases, including crystalline phases with $\mathbb{Z}_{N}$ symmetry ($N$ being an integer), which are characterized by periodic arrangements of Rydberg excitations.
A schematic of the phase diagram is given in Figure 11.
V.5.2 Phase diagram construction
In previous studies, the phase diagram of the Rydberg model has been explored using various techniques such as entanglement entropy [47], machine learning methods [48], and fidelity [49]. We obtain the phase diagram, with the first result presented in Figure 12. Here, we particularly focus on the $\mathbb{Z}_{2}$ and $\mathbb{Z}_{3}$ crystalline phases, which are typically observed in this model (See Figure 13 for a closer look at the $\mathbb{Z}_{3}$ crystalline phase). We considered a chain of length $L=50$, with long-range interactions truncated at the fourth nearest neighbor. The $1/r^{6}$ interaction term was approximated using an exponential form, as described in [50].
Figure 13: Detail of the phase diagram for the Rydberg Model around the $\mathbb{Z}_{3}$ crystalline phase. Due to the high resolution of the phase diagram, ripple-like features can be observed which signals a more complex phase to be determined.
Our results are consistent with the parameter ranges for each crystalline phase in the phase diagram, as reported in previous studies [50]. However, a noteworthy observation is that our method captures additional features that were not identified by earlier approaches. Specifically, we observe ripple-like features at high values of the parameters swept, which may indicate the presence of a more complex phase - a topic previously discussed in [51, 49, 52]. Additionally, we identify a line within each crystalline phase corresponding to the minimum of the reduced fidelity susceptibility, or equivalently, the maximum of the fidelity.
Notably, due to the high resolution of our phase diagram, we once again observe ripple-like features, consistent with previous findings in [49].
V.5.3 Order parameters
Next, we focus on the determination of an order parameter to detect each of these phase transitions. We initially start with a one-site RDM, to try to determine the different phases as seen in Figure 12. The projectors from the determined order parameter can be seen in Figure 14. As is evident that while it can determine the difference between the ordered and disordered phases, it fails to differentiate the different crystalline phases. Extending this to three sites results in the same projectors, which are unable to differentiate the different crystalline phases.
Figure 14: Rydberg model – The expectation for a 1-site observable that highlights the disordered phase versus the crystalline phases $\mathbb{Z}_{2}$ and $\mathbb{Z}_{3}$. Notably, the two crystalline phases cannot be distinguished using a 1-site observable.
It is not until a four-site RDM is used that the different disordered phases, $\mathbb{Z}_{3}$, $\mathbb{Z}_{2}$ can be characterized. This result is seen from the projectors of the observable in Figure 15.
Firstly, for the $\mathbb{Z}_{2}$ crystalline phase, the projector of the state |$\ket{0101}$ is used to detect this phase, showing that the agreed state of alternating spin up and spin down is evident in this phase, agreeing with previous literature.
Figure 15: Rydberg model – Highlights from the eigen decomposition of the order parameter on a 4-site RDM. The plots present the expectations of the projectors that reveal respectively, the crystalline (a) $\mathbb{Z}_{2}$ and (b) $\mathbb{Z}_{3}$ phases.
Similarly for the $\mathbb{Z}_{3}$ phase, the projector of the |$\ket{0001}$ was determined from the order parameter, which is also in agreement with previous literature. Expanding the RDM to include additional sites would make this symmetry more evident.
@@ Appendix A Construction of the phase diagram
for a fixed step $h$ (which defines the lattice of parameters) and $d_{i},d_{j}\in\{-1,0,1\}$. We can immediately verify the computational advantage of the approach since adjacent lattice points share a fidelity perturbation. For example $f(\boldsymbol{\lambda}_{i,j},h\mathbf{e}_{1})=f(\boldsymbol{\lambda}_{i,j+1},-h\mathbf{e}_{1})$. In Figure 16, the fidelity perturbations are represented with the symbol and placed between adjacent lattice points to emphasize the concept of sharing. By considering the finite differences approximation of the second derivative777 For a function $f:\mathbb{R}\to\mathbb{R}$, under sufficient smoothing conditions, we make use of the following approximation $\frac{\mathrm{d}^{2}f}{\mathrm{d}x^{2}}\approx$ $\frac{f(x+h)+f(x-h)-2f(x)}{h^{2}},$ (30) for some $h>0$. , and by noting that $f(\boldsymbol{\lambda}_{i,j},\mathbf{0})=1$ (for all valid $i,j$), we obtain
$\tilde{g}(\boldsymbol{\lambda}_{i,j})=$ $\frac{4-\sum_{\boldsymbol{\delta}\in\Omega}f(\boldsymbol{\lambda}_{i,j},\boldsymbol{\delta})}{h^{2}},$ (31)
@@ Appendix B Solution of the order parameter discovery problem
In this section, we expand on the solution of the order parameter discovery problem defined in (IV). Instead, the special case in (IV) is treated in Appendix C. We restate the problem for clarity and convenience, so
$\underset{M\in\mathcal{S}_{m}}{\min}$ $\sum_{(i,j)\in I^{+}\times I^{-}}\left(-\frac{\langle M\rangle_{i}^{2}}{p_{i}}+\frac{\langle M\rangle_{j}^{2}}{p_{j}}\right),$
@@
with $\langle M\rangle_{i}\coloneqq\operatorname{Tr}(\rho_{0}(\boldsymbol{\lambda}_{i})M)$ and $p_{i}\coloneqq\operatorname{Tr}(\rho_{0}(\boldsymbol{\lambda}_{i})^{2})$.
@@
Let $\rho_{i}=\rho_{0}(\boldsymbol{\lambda}_{i})$, and let $I$ be some finite index set. Assume $M\in\mathcal{S}_{m}$, then we have that
$\sum_{i\in I}\frac{\langle M\rangle_{i}^{2}}{p_{i}}=$ $\sum_{i\in I}\frac{\left(\operatorname{Tr}(\rho_{i}M)\right)^{2}}{p_{i}}$ (40a)
${=}$ $\sum_{i\in I}\frac{\left(\operatorname{\mathrm{vec}_{r}}(M)^{\dagger}\operatorname{\mathrm{vec}_{r}}(\rho_{i})\right)^{2}}{p_{i}}$ (40b)
$=$ $\mathbf{x}^{\dagger}\left(\sum_{i\in I}\frac{\mathbf{r}_{i}\mathbf{r}_{i}^{\dagger}}{p_{i}}\right)\mathbf{x},$ (40c)
with $\mathbf{x}\coloneqq\operatorname{\mathrm{vec}_{r}}(M)$ and $\mathbf{r}_{i}\coloneqq\operatorname{\mathrm{vec}_{r}}(\rho_{0}(\boldsymbol{\lambda}_{i}))$. We note that the vectors $\mathbf{x}$ and $\mathbf{r}_{i}$ belong to the set of vectorized Hermitians $\widehat{\mathcal{S}_{m}}$.
We use the result in (40a) to rewrite the optimization problem in (IV) in the equivalent [^8] Up to the positive constant (for the objective) $|I^{+}|\cdot|I^{-}|$. form
@@
$A\coloneqq$ $-\frac{1}{|I^{+}|}\sum_{i\in I^{+}}\frac{\mathbf{r}_{i}\mathbf{r}_{i}^{\dagger}}{p_{i}}+\frac{1}{|I^{-}|}\sum_{j\in I^{-}}\frac{\mathbf{r}_{j}\mathbf{r}_{j}^{\dagger}}{p_{j}}.$ (42)
This formulation confirms that the optimization is a quadratically constrained quadratic program as stated in Section IV. The matrix $A$ (determined by data) is Hermitian, so the objective in (41a) is real and well-defined (even in the absence of constraint (41c)). We expand on the condition for the non-degeneracy of the problem introduced in Section IV. We impose that the matrix $A$ is indefinite [61], where the Hermitian structure is fulfilled by the construction in (42). In other words, we require the matrix $A$ to have both (strictly) positive and negative eigenvalues999 As a simple example, this condition is not met when the sets $I^{+}$ and $I^{-}$ are identical. However, this would be incoherent since there is no distinction between the phases. .
@@
The optimization problem is non-convex since we do not assume $A\succcurlyeq 0$, indeed $A$ is required indefinite. However, as anticipated in Section IV, this optimization problem can be solved efficiently even in the case of the non-convexity of the objective (i.e., matrix $A$ is not PSD). Moreover, this is an exceptional case where strong duality101010 Strong duality is equivalent to the duality gap is zero, that is the difference between primal and dual solutions.  [28] holds, provided that Slater’s constraint qualification is fulfilled. That is, there exists an $\mathbf{x}\in\widehat{\mathcal{S}_{m}}$ such that the inequality constraint (41b) holds strictly (i.e. not tight). In our case, an example is $\mathbf{x}=\operatorname{\mathrm{vec}_{r}}(\mathbb{I}_{m})/\sqrt{m+\epsilon}$ for any $\epsilon>0$, so $\|x\|^{2}_{2}<1$.
@@
then the constraint is fulfilled when $\alpha\geq-\lambda_{\mathrm{min}}(A)$, so the dual optimal is $\alpha^{\star}=-\lambda_{\mathrm{min}}(A)>0$. The space of the primal solutions is given by the eigenspace corresponding to the minimum eigenvalue of $A-\lambda_{\mathrm{min}}(A)\mathbb{I}_{m^{2}}$, that is the null space of the latter. Given any matrix $A$, we denote the null space of $A$ (i.e., the set of solutions of the homogeneous equation $Ax=0$) by $\mathrm{Null}(A)$. Let
@@
with $\|x^{\star}\|^{2}_{2}=1$, be an optimal solution111111 We note that the vectorization operator $\operatorname{\mathrm{vec}_{r}}(\cdot):\mathbb{C}^{m\times m}\to\mathbb{C}^{m^{2}}$ is an isomorphism, so the expression $\mathbf{x}^{\star}=\operatorname{\mathrm{vec}_{r}}(M)$ implicitly means that the matrix $M$ is obtained uniquely from $\mathbf{x}^{\star}$ using the inverse of $\operatorname{\mathrm{vec}_{r}}$. , we show that if $\mathbf{x}^{\star}=\operatorname{\mathrm{vec}_{r}}(M)$, then $M=M^{\dagger}$. That is constraint (41c) is implied by the structure of matrix $A$.